# NeuRAD: Neural Rendering for Autonomous Driving

### v2

@@ 1 Introduction
Multiple works have applied NeRFs to automotive data [29, 18, 9, 32, 46, 44, 47, 34]. Neural Scene Graphs (NSG) [29] extend the original NeRF model [25] to dynamic automotive sequences by dividing the scene into a static background and a set of rigid dynamic actors with known location and extent, learning separate NeRFs for each. This enables editing the trajectories of both the ego-vehicle and all actors in the scene. The approach can be further improved by including semantic segmentation [18] or by using anti-aliased positional embeddings [46]. The latter enables NeRFs to reason about scale [3] which is essential for large-scale scenes. However, common for all these approaches is that they require many hours of training, limiting their applicability for scalable closed-loop simulation or data augmentation.
@@
More recent works [47, 44] rely on Instant NGP’s (iNGP) [27] learnable hash grids for embedding positional information, drastically reducing training and inference time. These methods are fast to train and generate realistic renderings in their respective settings, namely front-facing camera with ${}^{\circ}$ lidar. However, their performance in $360^{\circ}$ multicamera settings, which is common in many AD datsets [5, 43], is either unexplored [44] or is reported by the authors to be suboptimal [47]. Furthermore, both methods deploy simple lidar models and cannot model ray drop, a phenomenon important for closing the real-to-sim gap [21]. Lastly, using the iNGP positional embedding without anti-aliasing techniques limits performance, especially for larger scenes [4].
@@
In this paper, we present NeuRAD, an editable novel view synthesis (NVS) method. NeuRAD is designed to handle large-scale automotive scenes and to work well with multiple datasets out of the box. We find that modeling important sensor characteristics, such as rolling shutter, lidar ray dropping, and beam divergence, is essential for sensor-realistic renderings and learning accurate geometries. Nonetheless, our model features a simple network architecture, where static and dynamic elements are discerned only by their positional embeddings, making it a natural extension of recent methods to AD data. We verify NeuRAD’s generalizability and achieve state-of-the-art (SoTA) performance across five automotive datasets without any dataset-specific tuning.
@@
The contributions are summarized as follows. (1) Our method is the first to combine lidar sensor modeling with the ability to handle ${}^{\circ}$ camera rigs in a unified way, extending the applicability of NeRF-based methods for dynamic AD data. (2) We propose using a single network to model dynamic scenes, where dynamics and statics are separated only by their positional embeddings. (3) We propose simple, yet effective methods for modeling multiple key sensor characteristics such as rolling shutter, beam divergence, and ray dropping, and highlight their effect on performance. (4) Extensive evaluation using five popular AD datasets [1, 5, 43, 45, 10] shows that our method matches or outperforms state-of-the-art approaches across the board.
@@ 2 Related work
NeRFs for automotive data: Accurately simulating data for AD systems is a promising avenue for efficient testing and verification of self-driving vehicles. While game-engine-based methods [7, 31] have made a lot of progress, they struggle with scalable asset creation, real-to-sim gap, and diversity. NeRFs’ sensor-realistic renderings offer an attractive alternative, and consequently, multiple works have studied how to apply neural rendering techniques to automotive data. Neural Scene Graphs [29], Panoptic Neural Fields (PNF) [18] and Panoptic NeRF [9] all model the background and every actor as multi-layer perceptrons (MLPs), but struggle with large-scale scenes due to the MLPs limited expressiveness. S-NeRF [46] extends mip-NeRF 360 to automotive data similar to NSG by modeling each actor with a separate MLP, but requires day-long training, making it impractical for downstream applications. Block-NeRF [32] and SUDS [34] both focus on city-scale reconstruction. While handling impressive scale, Block-NeRF filters out dynamic objects and only models static backgrounds, and SUDS uses a single network for dynamic actors, removing the possibility of altering actor behavior.
@@
UniSim [47] is a neural sensor simulator, showcasing realistic renderings for PandaSet’s [45] front camera and ${}^{\circ}$ lidar. The method applies separate hash grid features [27] for modeling the sky, the static background, and each dynamic actor, and uses NSG-style [29] transformations for handling dynamics. For efficiency, the static background is only sampled near lidar points. Further, UniSim renders features from the neural field, rather than RGB, and uses a convolutional neural network (CNN) for upsampling the features and producing the final image. This allows them to reduce the number of sampled rays per image significantly. While efficient, multiple approximations lead to poor performance outside their evaluation protocol. In addition, the lidar occupancy has a limited vertical field of view and fails to capture tall, nearby structures which often becomes evident when using cameras with alternative mounting positions or wider lenses, e.g., nuScenes [5], Argoverse2 [43] or Zenseact Open Dataset (ZOD) [1]. In contrast, our method unifies static and sky modeling and relies on proposal sampling [4] for modeling occupancy anywhere. Further, UniSim’s upsampling CNN introduces severe aliasing and model inconsistencies, as camera rays must describe entire RGB patches whereas lidar rays are thin laser beams. In this work, we introduce a novel anti-aliasing strategy that improves performance, with minimal impact on runtime.
@@ 3.1 Scene representation and sensor modeling
$\mathbf{f}(\mathbf{r})=\sum^{N_{\mathbf{r}}}_{i=1}w_{i}\mathbf{f}_{i},\quad w_{i}=\alpha_{i}\prod^{i-1}_{j=1}(1-\alpha_{j}).$ (1)
Here, $\alpha_{i}$ represents the opacity at the point $\mathbf{x}_{i}=\mathbf{o}+\tau_{i}\mathbf{d}$, and $w_{i}$ the opacity times the accumulated transmittance along the ray up to $\mathbf{x}_{i}$. Inspired by its success in recovering high-quality geometry [28, 20], we represent the implicit geometry using a signed distance function (SDF) and approximate the opacity as
$\alpha_{i}=\frac{1}{1+e^{\beta s_{i}}},$ (2)
where $s_{i}$ is the SDF value at $\mathbf{x}_{i}$ and $\beta$ is a learnable parameter. While more accurate SDF formulations [35, 39] can provide better performance, they require gradient calculations for each 3D point, negatively impacting the runtime.
@@ 3.3 Automotive data modeling
[…]
@@ 3.4 Losses
Image losses: The image loss is computed patch-wise and summed over $N_{p}$ patches and consists of a reconstruction term and a perceptual term:
$\mathcal{L}^{\text{image}}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\lambda^{\text{rgb}}\mathcal{L}^{\text{rgb}}_{i}+\lambda^{\text{vgg}}\mathcal{L}^{\text{vgg}}_{i}.$ (4)
The reconstruction loss, $\mathcal{L}^{\text{rgb}}_{i}$, is the squared error between predicted and true pixel values. The perceptual loss, $\mathcal{L}^{\text{vgg}}_{i}$, is the same as proposed in pix2pixHD [38]. The two loss terms are weighted using the hyperparameters $\lambda^{\text{rgb}}$ and $\lambda^{\text{vgg}}$.
@@
$\mathcal{L}^{\text{lidar}}=\frac{1}{N}\sum_{i=1}^{N}(\lambda^{\text{d}}\mathcal{L}^{\text{d}}_{i}+\lambda^{\text{int}}\mathcal{L}^{\text{int}}_{i}+\lambda^{p_{d}}\mathcal{L}^{p_{d}}_{i}+\lambda^{\text{w}}\mathcal{L}^{\text{w}}_{i}),$ (5)
@@
$\mathcal{L}^{\text{w}}_{i}=\sum_{\tau_{i,j}>\epsilon}\left\|w_{ij}\right\|_{2},$ (6)
where $\tau_{i,j}$ is the distance from sample $\mathbf{x}_{ij}$ to the lidar observation for ray $i$. For dropped rays, weight decay is applied up until the specified sensor range. Lastly, we omit the commonly used eikonal loss, as it provided minimal benefits at a high computational cost.
@@ 3.5 Implementation details
NeuRAD is implemented in the collaborative, open-source project Nerfstudio [33]. We hope that our developed supporting structures such as data loaders and native lidar support will encourage further research into this area. We train our method for 20,000 iterations using the Adam [17] optimizer. Using a single Nvidia A100, training takes about 1 hour. See Appendix A for further details.
@@ 4 Experiments
To verify the robustness of our model, we evaluate its performance on several popular AD datasets: nuScenes [5], PandaSet [45], Argoverse 2 [43], KITTI [10], and ZOD [1]. To prove the robustness of our method we use the same model and hyperparameters on all datasets. We investigate novel view synthesis performance both for hold-out validation images and for sensor poses without any ground truth. Furthermore, we ablate important model components.
Table 1: Image novel view synthesis performance comparison to state-of-the-art methods across five datasets. *our reimplementation. ${}^{\dagger}$ modified baselines from [47, 44, 46]. ${}^{\mathsection}$ partial results due to training instability.
@@
NeuRAD (ours) 26.70 0.778 0.193
@@
NeuRAD (ours) 25.97 0.757 0.247
nuScenes ${}^{\dagger}$ [3, 46] 24.37 0.795 0.240
@@
KITTI MOT ${}^{\dagger}$ [34, 44] 23.12 0.821 0.135
@@
NeuRAD (ours) 26.01 0.779 0.090
Argo2 UniSim* ${}^{\mathsection}$ ${}^{\mathsection}$ ${}^{\mathsection}$
@@ 4.1 Datasets and baselines
Below, we introduce the datasets used for evaluation. The selected datasets cover various sensors, and the included sequences contain different seasons, lighting conditions, and driving conditions. Existing works typically use one or two datasets for evaluation and build models around assumptions about available supervision, limiting their applicability to new settings. Therefore, for each dataset, we compare our model to SoTA methods that have previously adopted said dataset, and follow their respective evaluation protocols. Similar to our method, UniSim [47] imposes few supervision assumptions, and we, therefore, reimplement the method and use it as a baseline for datasets where no prior work exists. See Appendix C for reimplementation details and Appendix B for further evaluation details.
PandaSet: We compare our method to UniSim [47] and an iNGP version with lidar depth supervision provided by UniSim. We use every other frame for training and the remaining ones for testing, and evaluate on the same 10 scenes as UniSim. We study two settings: one with lidar and one with front-facing camera (Panda FC) for direct comparison with the results reported in [47], and one with lidar and all six cameras capturing the full ${}^{\circ}$ field-of-view around the vehicle (Panda 360). We also evaluate UniSim on the full ${}^{\circ}$ setting using our reimplementation.
@@ 4.2 Novel view synthesis
Table 2: Lidar novel view synthesis performance comparison to state-of-the-art methods. Depth is median L2 error [$\mathrm{m}$]. Intensity is RMSE. Drop acc. denotes ray drop accuracy. Chamfer denotes chamfer distance, normalized with num. ground truth points [$\mathrm{m}$]. *our reimplementation.
@@
NeuRAD (ours) 0.01 0.076 96.4 3.1
Lidar: We measure the realism of our lidar simulation in terms of L2 median depth error, RMSE intensity error and ray drop accuracy. We complement the depth error with the Chamfer distance as it enables us to evaluate performance on dropped rays as well. We compare only to UniSim, evaluated on PandaSet, as no other baseline simulates point clouds. UniSim has no notion of ray dropping, hence we assume rays to be dropped past the reported lidar range. We see in Tab. 2 that NeuRAD decreases the depth error by an order of magnitude compared to UniSim in the front-camera setting. Our method generalizes well to the ${}^{\circ}$ setting, where similar results are reported. Furthermore, we show that NeuRAD is capable of simulating realistic point clouds, thanks to its high ray drop accuracy and low Chamfer distance. Fig. 4 further shows the importance of modeling ray drop effects for lidar simulation. As noted in the figure, lidar beams that hit the road far away tend to bounce off and not return. Similar effects occur for transparent surfaces, such as the car window illustrated in the figure, where the lidar beams shoot right through. Modeling these effects can increase the realism of simulated point clouds.
@@ 4.3 Novel scenario generation
In order for our method to be useful in practice, it must not only perform well when interpolating between views, but also when exploring new views, as examplified in Fig. 1. To that end, we investigate NeuRAD’s capability to generate images that are significantly different from those encountered during training. We adapt UniSim’s protocol on PandaSet, where the authors propose translating the ego vehicle sensors laterally two or three meters to simulate a lane shift, and extend the protocol to include one meter vertical shift, simulating other mounting positions. We further investigate “actor shift”, and rotate (±$\pm 0.5$ radians) or translate (±$\pm 2$ meters laterally) dynamic actors in the scene to simulate different actor behaviors. As no ground truth images exist, we report FID [12], with “no shift” for reference. The results in Tab. 3 show that NeuRAD is able to generalize to new viewpoints and learns meaningful actor representations. We also include results where we optimize the camera poses following [41], as this further increases sharpness.
Table 3: FID scores when shifting pose of ego vehicle or actors. *our reimplementation.

### v3

# supplementary

@@ 1 Introduction
Multiple works have applied NeRFs to automotive data [29, 18, 32, 46, 44, 47, 34]. Neural Scene Graphs [29] extend the original NeRF model [25] to dynamic automotive sequences by dividing the scene into static background and a set of rigid dynamic actors with known location and extent, learning separate NeRFs for each. This enables editing the trajectories of both the ego-vehicle and all actors in the scene. The approach can be further improved by including semantic segmentation [18] or by using anti-aliased positional embeddings [46]. The latter enables NeRFs to reason about scale [3] which is essential for large-scale scenes. However, common for all these approaches is that they require many hours of training, limiting their applicability for scalable closed-loop simulation or data augmentation.
@@
More recent works [47, 44] rely on Instant NGP’s (iNGP) [27] learnable hash grids for embedding positional information, drastically reducing training and inference time. Further, these methods generate realistic renderings in their respective settings, namely front-facing camera with 360 lidar. However, their performance in $360^{\circ}$ multicamera settings, which is common in many AD datsets [5, 43], is either unexplored [44] or is reported by the authors to be suboptimal [47]. Furthermore, both methods deploy simple lidar models and cannot model ray drop, a phenomenon important for closing the real-to-sim gap [21]. Lastly, using the iNGP positional embedding without anti-aliasing techniques limits performance, especially for larger scenes [4].
@@
In this paper, we present NeuRAD, an editable novel view synthesis (NVS) method, designed to handle large-scale automotive scenes and to work well with multiple datasets off the shelf. We find that modeling sensor characteristics, such as rolling shutter, lidar ray dropping, and beam divergence, is essential for sensor-realistic renderings. Further, our model features a simple network architecture, where static and dynamic elements are discerned only by their positional embeddings, making it a natural extension of recent methods to AD data. We verify NeuRAD’s generalizability and achieve state-of-the-art performance across five automotive datasets, with no dataset-specific tuning.
@@
Our contributions are as follows. (1) Our method is the first to combine lidar sensor modeling with the ability to handle 360 camera rigs in a unified way, extending the applicability of NeRF-based methods for dynamic AD data. (2) We propose using a single network to model dynamic scenes, where dynamics and statics are separated only by their positional embeddings. (3) We propose simple, yet effective methods for modeling multiple key sensor characteristics such as rolling shutter, beam divergence, and ray dropping, and highlight their effect on performance. (4) Extensive evaluation using five popular AD datasets shows that our method is state-of-the-art across the board.
@@ 2 Related work
NeRFs for automotive data: Accurately simulating data for AD systems is a promising avenue for efficient testing and verification of self-driving vehicles. While game-engine-based methods [7, 31] have made a lot of progress, they struggle with scalable asset creation, real-to-sim gap, and diversity. NeRFs’ sensor-realistic renderings offer an attractive alternative, and consequently, multiple works have studied how to apply neural rendering techniques to automotive data. NSG [29], Panoptic Neural Fields (PNF) [18] and Panoptic NeRF [9] all model the background and every actor as multi-layer perceptrons (MLPs), but struggle with large-scale scenes due to the MLPs limited expressiveness. S-NeRF [46] extends mip-NeRF 360 to automotive data similar to NSG by modeling each actor with a separate MLP, but requires day-long training, making it impractical for downstream applications. Block-NeRF [32] and SUDS [34] both focus on city-scale reconstruction. While handling impressive scale, Block-NeRF filters out dynamic objects and only models static backgrounds, and SUDS uses a single network for dynamic actors, removing the possibility of altering actor behavior.
@@
UniSim [47] is a neural sensor simulator, showcasing realistic renderings for PandaSet’s [45] front camera and 360 lidar. The method applies separate hash grid features [27] for modeling the sky, the static background, and each dynamic actor, and uses NSG-style [29] transformations for handling dynamics. For efficiency, the static background is only sampled near lidar points. Further, UniSim renders features from the neural field, rather than RGB, and uses a convolutional neural network (CNN) for upsampling the features and producing the final image. This allows them to reduce the number of sampled rays per image significantly. While efficient, multiple approximations lead to poor performance outside their evaluation protocol. In addition, the lidar occupancy has a limited vertical field of view and fails to capture tall, nearby structures which often becomes evident when using cameras with alternative mounting positions or wider lenses, e.g., nuScenes [5], Argoverse2 [43] or Zenseact Open Dataset (ZOD) [1]. In contrast, our method unifies static and sky modeling and relies on proposal sampling [4] for modeling occupancy anywhere. Further, UniSim’s upsampling CNN introduces severe aliasing and model inconsistencies, as camera rays must describe entire RGB patches whereas lidar rays are thin laser beams. In this work, we introduce a novel anti-aliasing strategy that improves performance, with minimal impact on runtime.
@@ 3.1 Scene representation and sensor modeling
$\vspace{-2mm}\mathbf{f}(\mathbf{r})=\sum^{N_{\mathbf{r}}}_{i=1}w_{i}\mathbf{f}_{i},\quad w_{i}=\alpha_{i}\prod^{i-1}_{j=1}(1-\alpha_{j}).$ (1)
Here, $\alpha_{i}$ represents the opacity at the point $\mathbf{x}_{i}=\mathbf{o}+\tau_{i}\mathbf{d}$, and $w_{i}$ the opacity times the accumulated transmittance along the ray up to $\mathbf{x}_{i}$. Inspired by its success in recovering high-quality geometry [28, 20], we represent the implicit geometry using a signed distance function (SDF) and approximate the opacity as $\alpha_{i}=1/({1+e^{\beta s_{i}}})$, where $s_{i}$ is the SDF value at $\mathbf{x}_{i}$ and $\beta$ is a learnable parameter. While more accurate SDF formulations [35, 39] can provide better performance, they require gradient calculations for each 3D point, negatively impacting the runtime.
@@ 3.3 Automotive data modeling
(a) original
(b) no modeling
(c) modeling lidar only
(d) modeling lidar + camera
@@ 3.4 Losses
Image losses: The image loss is computed patch-wise and summed over $N_{p}$ patches and consists of a reconstruction term $\mathcal{L}^{\text{rgb}}$ and a perceptual term $\mathcal{L}^{\text{vgg}}$:
$\vspace{-2mm}\mathcal{L}^{\text{image}}=\frac{1}{N_{p}}\sum_{i=1}^{N_{p}}\lambda^{\text{rgb}}\mathcal{L}^{\text{rgb}}_{i}+\lambda^{\text{vgg}}\mathcal{L}^{\text{vgg}}_{i}.\vspace{-1mm}$ (3)
The reconstruction loss is the squared error between predicted and true pixel values. The perceptual loss is the distance between VGG features for real and predicted patches [38]. $\lambda^{\text{rgb}}$ and $\lambda^{\text{vgg}}$ are weighting hyperparameters.
@@
$\vspace{-2.2mm}\mathcal{L}^{\text{lidar}}=\frac{1}{N}\sum_{i=1}^{N}(\lambda^{\text{d}}\mathcal{L}^{\text{d}}_{i}+\lambda^{\text{int}}\mathcal{L}^{\text{int}}_{i}+\lambda^{p_{d}}\mathcal{L}^{p_{d}}_{i}+\lambda^{\text{w}}\mathcal{L}^{\text{w}}_{i}),\vspace{-2.2mm}$ (4)
@@
$\mathcal{L}^{\text{w}}_{i}=\sum_{\tau_{i,j}>\epsilon}\left\|w_{ij}\right\|_{2},\vspace{-1.5mm}$ (5)
where $\tau_{i,j}$ is the distance from sample $\mathbf{x}_{ij}$ to the lidar observation for ray $i$. For dropped rays, weight decay is applied up until the specified sensor range. Noteably, we omit the commonly used eikonal loss, as it provided minimal benefits at a high computational cost.
@@ 3.5 Implementation details
NeuRAD is implemented in the collaborative, open-source project Nerfstudio [33]. We hope that our developed supporting structures such as data loaders and native lidar support will encourage further research into this area. We train our main method (NeuRAD) for 20,000 iterations using the Adam [17] optimizer. Using a single Nvidia A100, training takes about 1 hour. To showcase the scalability of our approach, we also design a larger model with longer training (NeuRAD-2x). See Appendix A for further details.
@@ 4 Experiments
To verify the robustness of our model, we evaluate its performance on several popular AD datasets: nuScenes [5], PandaSet [45], Argoverse 2 [43], KITTI [10], and ZOD [1]. To prove the robustness of our method we use the same model and hyperparameters on all datasets. We investigate novel view synthesis performance both for hold-out validation images and for sensor poses without any ground truth. Furthermore, we ablate important model components. More results, including a study on the real2sim gap as well as failure cases can be found in Appendix F and Appendix G.
Table 1: Image novel view synthesis performance comparison to state-of-the-art methods across five datasets. *our reimplementation. baselines from [47, 44, 46]. §partial results due to training instability. Bold/underline for best/second-best.
@@
NeuRAD (ours) 26.58 0.778 0.190
NeuRAD-2x (ours) 26.84 0.801 0.148
@@
NeuRAD (ours) 25.97 0.758 0.242
NeuRAD-2x (ours) 26.47 0.779 0.196
nuScenes Mip360 [3, 46] 24.37 0.795 0.240
@@
NeuRAD-2x (ours) 27.13 0.820 0.205
KITTI MOT SUDS [34, 44] 23.12 0.821 0.135
@@
NeuRAD (ours) 27.00 0.795 0.082
NeuRAD-2x (ours) 27.91 0.822 0.066
Argo2 UniSim* 23.22§ 0.661§ 0.412§
@@ 4.1 Datasets and baselines
Below, we introduce the datasets used for evaluation. The selected datasets cover various sensors, and the included sequences contain different seasons, lighting conditions, and driving conditions. Existing works typically use one or two datasets for evaluation and build models around assumptions about available supervision, limiting their applicability to new settings. Therefore, for each dataset, we compare our model to SoTA methods that have previously adopted said dataset, and follow their respective evaluation protocols. Similar to our method, UniSim [47] imposes few supervision assumptions, and we, therefore, reimplement the method (denoted Unisim) and use it as a baseline for datasets where no prior work exists. See Appendix C for reimplementation details and Appendix B for further evaluation details.
PandaSet: We compare our method to UniSim [47] and an iNGP version with lidar depth supervision provided by UniSim. We use every other frame for training and the remaining ones for testing, and evaluate on the same 10 scenes as UniSim. We study two settings: one with lidar and front-facing camera (Panda FC) for direct comparison with the results reported in [47], and one with lidar and all six cameras capturing the full 360 field-of-view around the vehicle (Panda 360). We also evaluate UniSim on the full 360 setting using our reimplementation.
@@ 4.2 Novel view synthesis
Table 2: Lidar novel view synthesis performance comparison to state-of-the-art methods. Depth is median L2 error [$\mathrm{m}$]. Intensity is RMSE. Drop acc. denotes ray drop accuracy. Chamfer denotes chamfer distance, normalized with num. ground truth points [$\mathrm{m}$].
@@
NeuRAD (ours) 0.01 0.061 96.1 1.9
Lidar: We measure the realism of our lidar simulation in terms of L2 median depth error, RMSE intensity error and ray drop accuracy. We complement the depth error with the Chamfer distance as it enables us to evaluate performance on dropped rays as well. We compare only to UniSim, evaluated on PandaSet, as no other baseline simulates point clouds. UniSim has no notion of ray dropping, hence we assume rays to be dropped past the reported lidar range. We see in Tab. 2 that NeuRAD decreases the depth error by an order of magnitude compared to UniSim in the front-camera setting. Our method generalizes well to the 360 setting, where similar results are reported. Furthermore, we show that NeuRAD is capable of simulating realistic point clouds, thanks to its high ray drop accuracy and low Chamfer distance. Fig. 4 further shows the importance of modeling ray drop effects for lidar simulation. As noted in the figure, lidar beams that hit the road far away tend to disperse and not return. Similar effects occur for transparent surfaces, such as the car window illustrated in the figure, where the lidar beams shoot right through. Modeling these effects can increase the realism of simulated point clouds.
@@ 4.3 Novel scenario generation
In order for our method to be useful in practice, it must not only perform well when interpolating between views, but also when exploring new views, as examplified in Fig. 1. To that end, we investigate NeuRAD’s capability to generate images from poses that are significantly different from those encountered during training. We adapt UniSim’s protocol on PandaSet, i.e., translating the ego vehicle sensors laterally two or three meters to simulate a lane shift, and extend the protocol to include one meter vertical shift, simulating other mounting positions. We further investigate “actor shift”, and rotate (±$\pm 0.5$ radians) or translate (±$\pm 2$ meters laterally) dynamic actors in the scene to simulate different actor behaviors. As no ground truth images exist, we report FID [12], with “no shift” for reference. The results in Tab. 3 show that NeuRAD is able to generalize to new viewpoints and learns meaningful actor representations. We also include results where we optimize the camera poses following [41], as this further increases sharpness.
Table 3: FID scores when shifting pose of ego vehicle or actors.